# Learning-based Distributed Model Predictive Control using Multi-Agent Bayesian Optimization

### v1

@@ IV-C Convergence Analysis of MABO
V Numerical Example
To examine the viability of the proposed learning-based distributed MPC scheme, we consider a heterogeneous multi-agent system where three linear systems with different dynamics must satisfy their local constraints and the coupling equality constraints (a desired distance between their first states) in a distributed manner. However, the local state constraints and the coupling constraints will be violated due to model-mismatch and disturbances. Let us consider three agents with the following dynamics
@@ V Numerical Example
$\boldsymbol{\mathrm{x}}_{3}^{k+1}=\begin{bmatrix}0.88&0.3\\ 0&1.1\end{bmatrix}\boldsymbol{\mathrm{x}}_{3}^{k}+\begin{bmatrix}0.0837\\ 0.21\end{bmatrix}u_{3}^{k}.$ (50c)
and choose an imperfect model for three local MPC schemes as
@@
where the disturbance $e_{1}^{k}$ is random, uncorrelated and uniformly distributed in the interval $\left[-0.1,0\right]$. Let us label the states of the agents as $\boldsymbol{\mathrm{x}}_{1}=\left[x_{1,1},x_{1,2}\right]^{\top}$, $\boldsymbol{\mathrm{x}}_{2}=\left[x_{2,1},x_{2,2}\right]^{\top}$ and $\boldsymbol{\mathrm{x}}_{3}=\left[x_{3,1},x_{3,2}\right]^{\top}$. We then consider the local constraints $0\leq x_{1,1}\leq 0.5$, $0\leq x_{2,1}\leq 2$ and $-2\leq x_{3,1}\leq 0$ on the first agent, second agent and third agent, respectively. The control input constraint $-0.5\leq u_{1,2,3}\leq 0.5$ is considered for all agents, and coupling constraints are defined as relative distances $d_{12}=1.5$, $d_{13}=1.5$ and $d_{23}=3$ ($d_{ij}=-d_{ji}$). The local cost function for each agent is then defined as
$L_{i}(\boldsymbol{\mathrm{x}}^{k},\boldsymbol{\mathrm{u}}_{i}^{k})=$ (52)
$10\left\|\boldsymbol{\mathrm{x}}_{i}^{c}\right\|_{2}^{2}+L_{i}\left(\boldsymbol{\mathrm{x}}_{i}^{k},\boldsymbol{\mathrm{u}}_{i}^{k}\right)+\boldsymbol{\mathrm{p}}^{\top}\cdot\text{max}\left(0,h_{i}\left(\boldsymbol{\mathrm{x}}_{i}^{k},\boldsymbol{\mathrm{u}}_{i}^{k}\right)\right),$
where
$\boldsymbol{\mathrm{x}}_{i}^{c}\in\mathbb{R}^{2}:=\left[x_{i,1}-x_{j,1}-d_{i,j},j\neq i,j=1,\ldots,m\right]^{\top}$. The local cost functions $L_{i}$ can be, for instance, a quadratic function, and the penalty vector is set to $\boldsymbol{\mathrm{p}}=\left[100,100\right]$.
@@
As observed in Fig. 1, the coupling constraints are not satisfied for a DMPC without learning as the desired distances shown in green color cannot be captured by the actual distances shown in cyan color. These coupling constraints can be satisfied by the proposed learning-based DMPC, shown in red color.
@@
Figure 1: The gray lines show the evolution of coupling constraints (desired distances) during the learning process. The green lines show the desired distances between the first states of agents. The cyan lines show the evolution of distances when a conventional DMPC without learning is used. The red lines show the evolution of distances using a learned DMPC.
@@
Fig. 2 shows the coupling states, which are the first states of the three agents. To make the situation more challenging for the first agent, we simultaneously set a constraint and a desired reference to zero so that the state $x_{1,1}$ must be as close as possible to 00 while avoiding the yellow unsafe zone. As observed, in the results obtained from the conventional DMPC, the first agent violates the constraint at zero due to the disturbance $e_{1}$ while this constraint violation is disappeared and the agent keeps its first state $x_{1,1}$ as close as possible to 00 using the proposed MABO-based DMPC.
@@
Figure 2: The gray lines show the evolution of the first states during the learning process. For the first state of the agent $1$, We consider the point 00 as reference and constraint simultaneously. The yellow region shows the unsafe zone. The cyan lines show the evolution of the first states when a conventional DMPC without learning is used. The red lines show the evolution of the first states using a learned DMPC.
@@
Figure 3: The gray lines show the evolution of control signals during the learning process. The cyan lines show the results when a conventional DMPC without learning is used while the red lines show the results obtained form a learned DMPC.
@@
As observed in Fig. 4, the proposed coordinated-based learning method using MABO outperforms the conventional BO without coordination in achieving the best closed-loop performances for each local MPC scheme. Consequently, we observe that the learned local MPC schemes using a BO without coordination cannot perfectly handle the coupling constraints, as shown in Fig. 5. Moreover, the BO-DMPC cannot guide the first state of the first agent $x_{1,1}$ close to the reference point 00 accurately, as shown in Fig. 6.
@@
Figure 4: The evolution of the local closed-loop performance of agents during the learning process. A comparative analysis between the proposed multi-agent BO (MABO) and conventional BO combined with DMPC.
@@
Figure 5: The evolution of coupling constraints. A comparative analysis between the proposed multi-agent BO (MABO) and conventional BO combined with DMPC.
@@
Figure 6: The evolution of the first states. A comparative analysis between the proposed multi-agent BO (MABO) and conventional BO combined with DMPC.
@@ VI Concluding Remarks
In this paper, we addressed a coordinated learning mechanism for distributed MPC scheme based on dual-composition where the local MPC models cannot capture the real multi-agent system accurately. The proposed method leveraged a multi-agent Bayesian optimization to provide a coordination framework for learning the parameterized DMPC aiming at improving the local closed-loop performances. We observed that each agent can refine its performance in a coordinated manner so that the global task of multi-agent system is satisfied.

### v2

@@ IV-C Convergence Analysis of MABO
Remark 5.
Convergence analysis of the non-convex ADMM is generally a challenging problem due to the absence of Fejér monotonicity [28]. One approach can be to instead prove that the non-convex EI function (39) is monotonically decreasing and then use Lemmas 1 and 2 to establish two properties of $\alpha_{\text{EI}}$ to facilitate the convergence analysis of the ADMM-based MABO. This is left as a topic for future work.
V Numerical Examples
To examine the viability of the proposed learning-based distributed MPC scheme, we consider two numerical examples: 1) heterogeneous linear multi-agent systems with simple dynamics, and 2) formation control problem for multiple Wheeled Mobile Robots (WMRs).
Example 1: Linear Multi-agent System
In this example, we consider three linear time-invariant (LTI) systems with different dynamics that must satisfy their local constraint and a coupled equality constraint (which is a desired distance between their first states) in a distributed manner. However, the local state constraints and the coupling constraints would be violated due to the model mismatch and disturbances. Let us consider three agents with the following dynamics
@@ V Numerical Examples
$\boldsymbol{\mathrm{x}}_{3}^{k+1}=\begin{bmatrix}0.88&0.3\\ 0&1.1\end{bmatrix}\boldsymbol{\mathrm{x}}_{3}^{k}+\begin{bmatrix}0.0837\\ 0.21\end{bmatrix}u_{3}^{k},$ (70c)
and choose an imperfect model for the local MPC schemes as
@@
where the disturbance $e_{1}^{k}$ is random, uncorrelated and uniformly distributed in the interval $\left[-0.1,0\right]$. Let us label the states of the agents as $\boldsymbol{\mathrm{x}}_{1}=\left[x_{1,1},x_{1,2}\right]^{\top}$, $\boldsymbol{\mathrm{x}}_{2}=\left[x_{2,1},x_{2,2}\right]^{\top}$ and $\boldsymbol{\mathrm{x}}_{3}=\left[x_{3,1},x_{3,2}\right]^{\top}$. We then consider the local constraints $0\leq x_{1,1}\leq 0.5$, $0\leq x_{2,1}\leq 2$ and $-2\leq x_{3,1}\leq 0$. The control input constraint $-0.5\leq u_{1,2,3}\leq 0.5$ is considered for all agents, and coupling constraints are defined as relative distances $d_{12}=1.5$, $d_{13}=1.5$ and $d_{23}=3$ (note that $d_{ij}=-d_{ji}$). The local cost function for each agent is defined as $L_{i}(\boldsymbol{\mathrm{x}}^{k},\boldsymbol{\mathrm{u}}_{i}^{k})=10\left\|\boldsymbol{\mathrm{x}}_{i}^{c}\right\|_{2}^{2}+l_{i}\left(\boldsymbol{\mathrm{x}}_{i}^{k},\boldsymbol{\mathrm{u}}_{i}^{k}\right)+\boldsymbol{\mathrm{p}}^{\top}\cdot\text{max}\left(0,h_{i}\left(\boldsymbol{\mathrm{x}}_{i}^{k},\boldsymbol{\mathrm{u}}_{i}^{k}\right)\right)$, where $\boldsymbol{\mathrm{x}}_{i}^{c}:=\left[x_{i,1}-x_{j,1}-d_{i,j}\right]_{j\neq i}\in\mathbb{R}^{m-1}$. The local cost function $l_{i}$ can be, for instance, a quadratic function, and the penalty vector is set to $\boldsymbol{\mathrm{p}}=\left[100,100\right]$.
@@
As observed from Fig. 2, the coupling constraints are not satisfied for a DMPC without learning as the desired distances shown in green color cannot be captured by the actual distances shown in cyan color. These coupling constraints can be however satisfied by invoking the proposed learning-based DMPC, shown in red color.
@@
Figure 2: The gray lines show the evolution of the actual distances during the learning process. The green lines show the desired distances between the first states of agents. The cyan lines show the evolution of the actual distances when a conventional DMPC without learning is used. The red lines show the evolution of the actual distances using a learned DMPC.
@@
Fig. 3 shows the coupling states, which are the first states of the three agents. To make the situation more challenging for the first agent, we simultaneously set a constraint and a desired reference to zero so that the state $x_{1,1}$ must be as close as possible to 00 while avoiding the yellow unsafe zone. As observed, in the results obtained from the conventional DMPC, the first agent violates the constraint at zero due to the disturbance $e_{1}$ while this constraint violation disappears, and the agent keeps its first state $x_{1,1}$ close to 00 using the proposed MABO-based DMPC. The three control inputs are shown in Fig. 4.
@@
Figure 3: The gray lines show the evolution of the first states during the learning process. For the first state of the agent $1$, we consider the point 00 as reference and constraint simultaneously. The yellow region shows the unsafe zone. The cyan lines show the evolution of the first states when a conventional DMPC without learning is used. The red lines show the evolution of the first states using the proposed learned DMPC scheme.
@@
Figure 4: The gray lines show the evolution of control signals during the learning process. The cyan lines show the results when a conventional DMPC without learning is used while the red lines show the results obtained from using the learned DMPC.
@@
As observed in Fig. 5, the proposed coordinated learning method using MABO outperforms conventional BO without coordination in achieving the best closed-loop performance for each local MPC scheme. Consequently, we observe that the learned local MPC schemes using BO without coordination cannot perfectly handle the coupling constraints, as shown in Fig. 6. Moreover, the BO-DMPC cannot drive the first state of the first agent $x_{1,1}$ to the reference point 00 accurately, as shown in Fig. 7.
@@
Figure 5: The evolution of the local closed-loop performance of agents during the learning process is shown comparing the proposed multi-agent BO (MABO) and conventional BO combined with DMPC.
@@
Figure 6: The evolution of the coupling constraints is shown here comparing the proposed MABO and conventional BO combined with DMPC.
@@
Figure 7: The evolution of the first states is shown here comparing the proposed MABO and conventional BO combined with DMPC.
Example 2: Formation Control
We define the WMR model as
$\boldsymbol{\mathrm{f}}(\boldsymbol{\mathrm{x}},\boldsymbol{\mathrm{u}})=\begin{bmatrix}\cos(\psi)&0\\ \sin(\psi)&0\\ 0&1\end{bmatrix}\boldsymbol{\mathrm{u}},$ (72)
where $\boldsymbol{\mathrm{x}}=\left[x,y,\psi\right]^{\top}$ and $\boldsymbol{\mathrm{u}}=\left[v,\omega\right]^{\top}$ are the state and control input vectors, respectively. The position coordinates of the WMR are labeled $x,y$, and $\psi$ is the angle of orientation of the robot. The control inputs $v$ and $\omega$ are the linear and angular velocities, respectively. To discretize the above continuous model, we use a fourth-order Runge-Kutta (RK4) integrator that provides the discretized function $\boldsymbol{\mathrm{f}}_{d}$ of the WMR model as $\boldsymbol{\mathrm{x}}(k+1)=\boldsymbol{\mathrm{f}}_{d}\left(\boldsymbol{\mathrm{x}}(k),\boldsymbol{\mathrm{u}}(k)\right)$. We use two imperfect models of the real system for the second and third local MPC schemes. These two models are affected by the constant uncertainty in their control inputs such that we have the two imperfect models of (72) as $\boldsymbol{\mathrm{f}}_{2}\left(\boldsymbol{\mathrm{x}}_{2},d_{2}\boldsymbol{\mathrm{u}}_{2}\right)$ and $\boldsymbol{\mathrm{f}}_{3}\left(\boldsymbol{\mathrm{x}}_{3},d_{3}\boldsymbol{\mathrm{u}}_{3}\right)$. We select $d_{2}=0.2,d_{3}=1.5$ and consider that three WMRs are required to achieve a triangular formation. As observed in Fig. 8, the DMPC scheme (28) without learning cannot achieve the target triangular formation due to the uncertainties in the models that second and third WMRs employ. We then learn the parametric local MPC cost functions using the proposed MABO in order to approach the target formation as shown in Fig. 9. As observed in Fig. 10, we run the proposed MABO-DMPC for $360$ episodes to achieve the best local closed-loop performance leading to the desired triangular formation.
Figure 8: Formation control using the proposed DMPC scheme (28) with wrong MPC models $\boldsymbol{\mathrm{f}}_{2}$ and $\boldsymbol{\mathrm{f}}_{3}$. The red triangle shows the target formation while the cyan triangle is the actual formation. The purple, green, and blue lines show the trajectories of WMR 1, WMR 2, and WMR 3, respectively.
Figure 9: Formation control using the proposed learning-based parametric DMPC scheme (33) with wrong MPC models $\boldsymbol{\mathrm{f}}_{2}$ and $\boldsymbol{\mathrm{f}}_{3}$. The red triangle shows the target formation while the cyan triangle is the actual formation achieved after learning. The purple, green, and blue lines show the trajectories of WMR 1, WMR 2, and WMR 3, respectively. The gray lines show the evolution of the triangular formation during the MABO-based learning process.
Figure 10: The evolution of the local closed-loop performance (cost) of WMRs during the learning process.
@@ VI Concluding Remarks
In this paper, we developed a new coordinated learning framework for the design of distributed MPC schemes (DMPC) invoking dual composition assuming that the local MPC models are unable to capture the real multi-agent system accurately. The proposed method leveraged the machinery of multi-agent Bayesian optimization leading to a coordinated framework for learning the parameterized DMPC aiming at improving the local closed-loop performance. We observed significant improvements in allowing each agent to refine its performance in a coordinated manner so that the global task the multi-agent system is assigned to is satisfied.